# Gazing Into Missteps: Leveraging Eye-Gaze for Unsupervised Mistake Detection in Egocentric Videos of Skilled Human Activities

### v4

@@ 2 Related Work
Use of gaze in egocentric vision While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications. Specifically, previous investigations focused on discovering object usage [3], detecting privacy-sensitive situations [38], finding attended objects [27], assisting large language models in classification tasks [16], supporting radiologists to collect segmentation masks by looking at regions of interest during image diagnosis [42], improving egocentric human motion prediction [45], and aiding natural language processing tasks [36]. We show the effectiveness of gaze in mistake detection. Our method compares gaze trajectories predicted from visual data with gaze estimated through a gaze tracker to identify mistakes when predictions deviates from the ground truth.
@@ 4.4 Comparison with the state of the art
EPIC-Tent Table 2 compares the proposed mistake detection approach with competitors on EPIC-Tent, according to the three considered levels of supervision. C2F outperforms TimeSformer in all evaluated metrics, particularly in the F1 score ($0.58$ of C2F vs $0.49$ of TimeSformer and $0.36$ of the random baseline), suggesting that C2F is more adept at capturing temporal reasoning, which is crucial for identifying mistakes in dynamic activities. However, the reliance on fully labeled datasets poses limitations for both methods.
@@
One-class methods for anomaly detection adapted to take only gaze as input, namely, TrajREC (G) and MoCoDAD (G) show minor improvements over the random baseline in terms of AUC score ($0.50-0.51$ vs $0.51$) and only small improvements in F1 score ($0.40-0.43$ vs $0.36$). High recall values, paired with low precision, suggest that such methods tend to classify most clips as mistakes. Incorporating hand skeleton data instead of gaze, namely, TrajREC (H) and MoCoDAD (H), leads to a slight improvement, as evidenced by the F1 scores and AUC values of $0.44$ and $0.45$ respectively, with AUC values of $0.55$ and $0.60$ However, treating gaze as an additional keypoint, beside hand key-points (G+H configurations) yields lower F1 scores (0.42 for TrajREC (H+G) and 0.43 for MoCoDAD (H+G)) and AUC values (0.53 and 0.56), possibly be due to the fact that gaze trajectories do not always align with hand movements. Combining (H) and (G) models through late fusion (H+G configurations denoted with ) improves the results, with MoCoDAD (H+G) achieving an F1 score of 0.49 and an AUC of 0.65, which suggests that the signals captured by gaze-based analysis and hand-based analysis are complementary. The one-class approach based on GLC [19] gaze prediction shows improved results compared to previous one-class methods being only slightly less effective than the late fused method, despite not analyzing any hand-based information. This highlights the soundness of leveraging gaze analysis for mistake detection. Finally, the proposed method based on gaze completion obtains best results yielding an AUC of $0.69$, an F1-score of $0.52$, a precision of $0.37$, and a recall of $0.85$, which amount to relative improvements of $+5/13\%$ [^6] We compute the relative improvement of $b$ with respect to $a$ as $\frac{b-a}{b}$ with respect to the best approach GLC and $+35/44\%$ with respect to the random baseline. Late-fusing our approach with MoCoDAD (H) achieves enhanced results with an AUC of $0.72$ and an F1 score of $0.54$, which improve over compared approaches, suggesting that gaze analysis can further benefit from integration with approaches based on different cues. It is worth noting that our best method remarkably achieves the same AUC score of $0.72$ of the best supervised approach and a comparable $F1$ score ($0.54$ vs $0.58$) without access to labels at training.
@@
We compare unsupervised approaches in the bottom part of Table 2. Similarly to the one-class case, TrajREC (H) and MoCoDAD (H) slightly improve over the random baseline (e.g., $0.40$ and $0.41$ vs $0.36$ of the random baseline in F1 score). GLC outperforms these former two methods obtaining an F1 score of $0.44$ and an AUC score of $0.61$, which are lower than the scores of $0.46$ and $0.66$ obtained in the one-class setting. In these settings, the proposed method achieves an F1 score of $0.51$, which is comparable to the score obtained in one-class settings $0.52$ with similar and recall values and the same AUC of $0.69$, despite the unsupervised setting being more challenging. Late fusion with MoCoDAD (H) brings some additional improvements, with an F1 score of 0.52 and AUC of 0.70.
@@
Table 3 compares the proposed method with competitors on HoloAssist, which presents a more varied and expansive context compared to EPIC-Tent, including different sorts of activities, which makes it significantly more challenging to perform mistake detection. This is evidenced by the fact that the random baseline achieves an F1-score of only $0.04$. One-class baselines like TrajREC and MoCoDAD bring improvements over the baseline but tend to overly classify clips as mistakes (Recall $0.96$ and $0.94$).In line with the trends observed in the EPIC-Tent dataset, the results indicate that utilizing hand keypoints leads to improvements in this context. Specifically, TrajREC (H) slightly outperforms MoCoDAD (H) within the HoloAssist framework. GLC demonstrates enhancements in AUC and F1 scores, along with more balanced precision and recall metrics. Our methodology achieves an AUC of $0.61$ and an F1 score of $0.22$ in one-class settings, with the best result obtained through the integration of MoCoDAD using late fusion (F1 of $0.26$ and AUC of $0.63$). Performance is lower in the unsupervised scenario. TrajREC (G) and MoCoDAD (G) demonstrate limitations in effectiveness, while the (H) approaches perform slightly better. Nevertheless, our approach achieves an AUC score of 0.59 and an F1 score of 0.18, highlighting the robustness of our method across evaluation settings and showing improvements over GLC, with relative gains of $+9\%$ and $+80\%$, respectively. Combining with MoCoDAD (H) further boosts performance.
Results on IndustReal reported in Table 4 confirm previous observations, with TrajREC and MoCoDAD bringing small improvements over the random baseline in H, G and H+G configurations, and our method outperforming competitors and improving over GLC by systematic margins both in one-class ($+5\%$ in AUC and $+14\%$ in F1) as well as in unsupervised settings ($+6\%$ in AUC score and comparable F1 performance), while late fusion with MoCoDAD (H) does not boost performance due to the reduced performance of MoCoDAD in this scenario.

### v5

@@ 2 Related Work
Use of gaze in egocentric vision While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications. Specifically, previous investigations focused on discovering object usage [3], detecting privacy-sensitive situations [38], finding attended objects [27], assisting large language models in classification tasks [16], enhancing visual tasks [46, 44], improving egocentric human motion prediction [45], and aiding natural language processing tasks [36]. We show the effectiveness of gaze in mistake detection. Our method compares gaze trajectories predicted from visual data with gaze estimated through a gaze tracker to identify mistakes when predictions deviates from the ground truth.
@@ 4.4 Comparison with the state of the art
EPIC-Tent Table 2 compares the proposed mistake detection approach with competitors on EPIC-Tent, according to the three considered levels of supervision. C2F outperforms TimeSformer in all evaluated metrics, particularly in the F1 score ($0.58$ of C2F vs $0.49$ of TimeSformer and $0.36$ of the random baseline), suggesting that C2F is more adept at capturing temporal reasoning, which is crucial for identifying mistakes in dynamic activities. However, the reliance on fully labeled datasets poses limitations for both methods. One-class methods for anomaly detection adapted to take only gaze as input, namely, TrajREC (G) and MoCoDAD (G), show minor improvements over the random baseline in terms of AUC score ($0.50-0.51$ vs $0.51$) and only small improvements in F1 score ($0.40-0.43$ vs $0.36$). High recall values, paired with low precision, suggest that these methods tend to classify most clips as mistakes. Incorporating hand skeleton data instead of gaze, namely, TrajREC (H) and MoCoDAD (H), leads to slight improvements, as evidenced by F1 scores of $0.44$ and $0.45$, and AUC values of $0.55$ and $0.60$. Combining (H) and (G) models through late fusion (denoted with ) improves the results, with MoCoDAD (H+G) achieving an F1 score of 0.49 and an AUC of 0.65, suggesting that the signals captured by gaze-based and hand-based analyses are complementary. The one-class approach based on GLC [19] gaze prediction shows improved results compared to previous one-class methods, being only slightly less effective than the late fused method, despite not analyzing any hand-based information. This highlights the value of leveraging gaze analysis for mistake detection. “‘
@@
Finally, the proposed method based on gaze completion obtains the best results, yielding an AUC of $0.69$, an F1-score of $0.52$, a precision of $0.37$, and a recall of $0.85$, which amount to relative improvements of $+5/13\%$ [^6] We compute the relative improvement of $b$ with respect to $a$ as $\frac{b-a}{b}$ with respect to the best approach GLC and $+35/44\%$ with respect to the random baseline. Late-fusing our approach with MoCoDAD (H) achieves enhanced results with an AUC of $0.72$ and an F1 score of $0.54$, improving over compared approaches, suggesting that gaze analysis can further benefit from integration with approaches based on different cues. It is worth noting that our best method remarkably achieves the same AUC score of $0.72$ as the best supervised approach and a comparable $F1$ score ($0.54$ vs $0.58$) without access to labels during training. We compare unsupervised approaches in the bottom part of Table 2. Similarly to the one-class case, TrajREC (H) and MoCoDAD (H) slightly improve over the random baseline (e.g., $0.40$ and $0.41$ vs $0.36$ of the random baseline in F1 score). GLC outperforms these former two methods obtaining an F1 score of $0.44$ and an AUC score of $0.61$, which are lower than the scores of $0.46$ and $0.66$ obtained in the one-class setting. In these settings, the proposed method achieves an F1 score of $0.51$, which is comparable to the score obtained in one-class settings $0.52$ with similar and recall values and the same AUC of $0.69$, despite the unsupervised setting being more challenging. Late fusion with MoCoDAD (H) brings some additional improvements, with an F1 score of 0.52 and AUC of 0.70.
@@
Table 3 compares the proposed method with competitors on HoloAssist, which presents a more varied and expansive context than EPIC-Tent, making mistake detection more challenging. The random baseline achieves an F1 score of only $0.04$. One-class methods like TrajREC and MoCoDAD improve over the baseline but tend to classify most clips as mistakes, with high recall values ($0.96$ and $0.94$). Hand keypoint-based methods show improvements, with TrajREC (H) slightly outperforming MoCoDAD (H). GLC demonstrates better AUC and F1 scores, with more balanced precision and recall metrics. Our approach achieves an AUC of $0.61$ and an F1 score of $0.22$ in one-class settings, with the best results from late fusion with MoCoDAD, yielding an F1 score of $0.26$ and an AUC of $0.63$. In the unsupervised scenario, TrajREC (G) and MoCoDAD (G) show limited effectiveness, while the (H) approaches perform slightly better. Our method achieves an AUC of $0.59$ and an F1 score of $0.18$, showing robustness across evaluation settings and relative improvements over GLC, with gains of $+9\%$ and $+80\%$, respectively. Combining with MoCoDAD (H) further enhances performance.
Results on IndustReal, shown in Table 4, confirm the trends observed in HoloAssist. TrajREC and MoCoDAD bring small improvements over the random baseline in H, G, and H+G configurations. Our method outperforms competitors with improvements over GLC of $+5\%$ in AUC and $+14\%$ in F1 in one-class settings, and $+6\%$ in AUC in unsupervised settings, while late fusion with MoCoDAD (H) does not improve performance due to MoCoDAD’s reduced effectiveness in this scenario.
@@ 5 Conclusion
Acknowledgments
This research has been supported by the project Future Artificial Intelligence Research (FAIR) – PNRR MUR Cod. PE0000013 - CUP: E63C22001940006.
Supplementary Material
Appendix A Implementation details
Our model implementation is primarily based on the approach outlined in [19], with hyperparameters adjusted accordingly. Below, we elaborate on the key aspects of our implementation, highlighting the differences and specific choices made to enhance the model’s performance.
Code Availability: To ensure reproducibility and provide further implementation details, we will share the complete codebase upon publication.
Figure 5: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake. (a) Entropy. (b) Euclidean distance between the two trajectories. (c) Dynamic Time Warping (DTW). (d) Average value of the ground truth trajectory at the predicted heatmaps.
Stride Adjustment: In contrast to the stride of 8 used in [19], we opted for a stride of 1 during training. This modification allows the model to process consecutive frames without skipping, leading to finer granularity in temporal feature extraction. Our experiments indicated that this adjustment results in marginal improvements in both gaze estimation and mistake prediction accuracy.
Overfitting Prevention: To mitigate the risk of overfitting, we incorporated a weight decay parameter set to 0.07. This regularization technique helps in controlling the complexity of the model by penalizing large weights, thereby promoting generalization to unseen data.
[…]
Training Loss Following [19], we consider gaze prediction as defining a probability distribution over the 2D image plane of each input frame. Our proposed method leverages an architecture modified for gaze completion to predict missing segments of gaze trajectories. Ablations on single frames showed that looking at sequences of frames, which create a trajectory, is more effective for detecting anomalies due to the importance of changes over time. We train the model by minimizing the sum of the Kullback–Leibler divergence between the predicted gaze maps $\hat{P}(i)$ and the ground truth ones $Q(i)$ at each frame $i$:
$L_{\text{KL}}(\hat{P}\parallel Q)=\sum_{i}\hat{P}(i)\log\left(\frac{\hat{P}(i)}{Q(i)}\right)$ (6)
Graphical Illustration of scoring functions Figure 5 illustrates the scoring function considered in this study. Entropy is the only scoring function which does not require any ground truth gaze as input, but only evaluates the level of uncertainty of the predicted heatmaps. The Euclidean and DTW scoring functions compute two forms of distances between predicted and ground truth trajectories. The heatmap scoring function evaluates the probability of predicted gaze under the points indicated by the ground truth trajectory. The heatmap scoring function achieves best results in our experiments. The main paper reports the formal definition of such scoring functions.
MoCoDAD Baseline
Following methodologies from [8], we employ a sliding window approach to segment each agent’s gaze/hands history. A window size of 8 frames is utilized, with the initial 4 frames dedicated to condition setting and the subsequent frames for the diffusion process. Hyperparameters are set as $\lambda_{1}=\lambda_{2}=1$. Training proceeds end-to-end using the Adam optimizer with a learning rate of $1\times 10^{-4}$, employing exponential decay over 25 epochs. The diffusion process utilizes $\beta_{1}=1\times 10^{-4}$, $\beta_{T}=2\times 10^{-2}$ for $T=10$, and incorporates the cosine variance scheduler.
TrajREC Baseline
We followed the implementation proposed in TrajREC [39] official code release [^10] https://github.com/alexandrosstergiou/TrajREC, adapting it for gaze/hands trajectory analysis. The approach encodes temporally occluded gaze/hands trajectories, jointly learns latent representations of occluded segments, and reconstructs trajectories based on expected motions across different temporal segments.
For both methods, if a frame does not contain gaze or hand keypoints, we exclude that frame from the score calculation for the segment.
Method Fusion F1 Recall Precision
Gaze Prediction // 0.37 0.65 0.26
Gaze Completion CH 0.38 0.67 0.29
Gaze Completion CH + CORR 0.40 0.70 0.31
Table 5: Comparison of GLC and the proposed Gaze Completion approach for Gaze Estimation on EPIC-Tent.
Action Type Classification
To assess whether the type of performed action affects the performance of our method, we grouped actions contained in all three datasets into four categories: Hand-Eye Coordination, Object Manipulation, Task Preparation, and Inspection/Verification. For categorization, we prompted GPT-4 with a full list of actions using the following prompt:
In the context of how gaze affects actions, organize the following actions into groups that align with gaze literature. Group the actions from those that involve the most fine-grained gaze coordination to those that involve less gaze precision.
The list was then manually revised. The full classification is shown in Table 6
Category Dataset Actions
Hand-Eye Coordination EpicTent assemble, insert stake, insert support, insert support tab, tie top
HoloAssist touch, place, lift, press, flip, unscrew, rotate, slide, insert, close, turn, screw, disassemble
IndustReal fit, plug, tighten, loosen
Object Manipulation EpicTent spread tent, place guyline
HoloAssist adjust, empty, drop, clean, make, pour, split, mix-stir, stack-pile, load, mount, lock, unlock, shift, grab, pull
IndustReal put, take, pull
Task Preparation EpicTent pickup/open stakebag, pickup/open supportbag, pickup/open tentbag
HoloAssist withdraw, exchange, hold, break, approach, stand, align
IndustReal align
Inspection/Verification EpicTent instruction, place ventcover
HoloAssist inspect, validate, point, tap, click, push
IndustReal check, browse
Table 6: Classification of actions by category across datasets based on gaze involvement.
Method Sup. Level Overall F1 Overall AUC Hand-Eye Coord. Object Manip. Task Prep. Inspect/Verif.
F1 AUC F1 AUC F1 AUC F1 AUC
Random // 0.36 0.51
TimeSformer [2] Fully Supervised 0.49 0.67 0.452 0.615 0.474 0.636 0.551 0.691 0.532 0.678
C2F [35] Fully Supervised 0.58 0.72 0.506 0.600 0.5622 0.771 0.5138 0.686 0.741 0.857
GLC [19] One-Class 0.46 0.66 0.524 0.704 0.495 0.665 0.425 0.579 0.396 0.556
Ours One-Class 0.52 0.69 0.741 0.839 0.612 0.734 0.489 0.643 0.244 0.543
Ours + MoCoDAD (H)* One-Class 0.54 0.72 0.753 0.872 0.631 0.764 0.498 0.657 0.257 0.543
GLC [19] Unsupervised 0.44 0.61 0.542 0.694 0.474 0.657 0.406 0.563 0.338 0.526
Ours Unsupervised 0.51 0.69 0.711 0.839 0.603 0.723 0.483 0.637 0.240 0.531
Ours + MoCoDAD (H)* Unsupervised 0.52 0.70 0.714 0.862 0.602 0.754 0.489 0.646 0.253 0.535
    Late fusion
Table 7: Mistake detection results on EPIC-Tent by category. Best results are in bold, second best results are underlined.
Baseline Future frames F1 Precision Recall
Gaze Completion 1 0.46 0.39 0.59
Gaze Completion 2 0.47 0.38 0.62
Gaze Completion 3 0.47 0.37 0.63
Gaze Completion 4 0.49 0.34 0.88
Table 8: Performance ablation for different prediction lengths. Smaller windows yield higher precision but lower recall. The best F1 score is achieved when predicting 4 frames into the future.
Appendix B Additional Ablations
This section reports additional ablations which could not be included in the submitted paper due to space limits.
B.1 Performance Comparison Across Action Types
Table 7 compares the performance of the considered baselines and our proposed method across different action types.
The fully supervised C2F method achieves overall F1 and AUC scores of 0.58 and 0.72, respectively, maintaining stable performance across the four action types. The best performance is observed in Inspect/Verify actions (F1: 0.74, AUC: 0.85), likely due to the strong visual cues inherent to these tasks (e.g., instruction sheets).
In contrast, under both One-Class and Unsupervised scenarios, the proposed gaze-based approaches show varying performance depending on the action type. Stronger results are observed in tasks requiring Hand-Eye Coordination and Object Manipulation skills. Under the One-Class supervision level, our method achieves an F1 score of 0.741 and an AUC of 0.839 for hand-eye coordination tasks (+21% vs Overall AUC). This indicates its effectiveness in learning “normal” attention patterns and detecting mistakes during complex actions where gaze and motor coordination are crucial.
Conversely, for simpler actions, such as Task Preparation and Inspect/Verify, the proposed approaches are less effective. This is likely due to the high gaze variability inherent in less skill-intensive tasks. For instance, under the One-Class supervision level, our method achieves an F1 score of 0.257 and an AUC of 0.543 (-24% vs Overall AUC) for task preparation actions.
B.1.1 Performance of the proposed gaze completion approach vs standard gaze prediction
Results in main paper Table 1 compared the performance of the proposed mistake detection method based on gaze completion versus different methods, including a baseline method based on the standard gaze prediction task implemented with the method of [19]. In Table 5, we instead compare the performance of the proposed gaze completion approach with standard gaze prediction based on [19] on the EPIC-Tent dataset. Just using channel fusion brings a performance boost, achieving an F1-score of $0.38$, recall of $0.67$, and precision of $0.29$, while combining channel and correlation fusion brings best results with an F1-score of $0.40$, recall of $0.70$, and precision of $0.31$, suggesting that conditioning on partial trajectories makes gaze prediction less uncertain and the proposed approach can leverage the informative content provided by the input trajectory surpassing the performance of standard gaze prediction. Moreover, the performance in Table 5 correlates with the results in Table 1 of the main paper, suggesting that accurate gaze prediction enhances mistake detection performance. Specifically, the proposed approach excels in gaze prediction for “Correct execution” frames, although it loses accuracy for “Mistake” frames. Given that “Correct execution” frames are generally more frequent, the F1 score improves overall, but the gap in prediction accuracy between “Correct execution” and “Mistake” frames widens. This discrepancy, however, benefits trajectory-based comparisons in mistake detection, as the increased accuracy in “Correct execution” frames helps to better identify errors in subsequent frames.
B.2 Length of prediction and performance
Table 8 ablates performance for different prediction lengths. Smaller windows lead to higher precision due to short future trajectories being more predictable, but also lower recall, with the best F1 score when predicting 4 frames into the future.
As the prediction window extends from 1 to 4 frames, the model’s recall improves, indicating more mistakes detected. However, this is offset by a reduction in precision, leading to more false positives.
Figure 6: Length of prediction.
B.3 Chosen thresholds and sensitivity
We report F1 scores obtained at each method’s optimal thresholds, which we’ll report in the paper. Figure 6 shows how the F1 score of our best method (Unsupervised - Ours, Table 2 of main paper) changes when varying the threshold. Performance is stable for a range of threshold values.
(a) Correct prediction of Correct action.
(b) Correct prediction of Mistake Action.
Figure 7: Qualitative examples. The first four columns represent the inputs (with the input gaze 2D points highlighted in orange). The latter four columns show the predicted outputs in the form of heatmaps.
B.4 Qualitative Results and Failure Cases
Figure 7 illustrates the performance of various baselines and the proposed approach for both correct predictions in Correct Execution cases (a) and in Incorrect Execution cases (b). The top row displays the ground truth, followed by predictions from the GLC method, our proposed “Channel” approach, and, at the bottom, the ”Gaze Frame Correlation” approach. The last four columns display the predicted heatmap, where red peaks symbolize the 2D gaze predicted points.
Figure 7(a) focuses on correct predictions related to Correct Execution. The first row shows the actual gaze coordinates. Notably, in the second row corresponding to GLC, the predicted heatmaps exhibit inconsistencies, with varying peaks across consecutive frames. In contrast, our proposed method leverages temporal information to produce temporally consistent predictions. The “Channel” approach demonstrates better consistency than GLC, while the “Gaze Frame Correlation” method generates more defined heatmaps with fewer, more localized peaks around the gaze region. In this case, a Correct Execution is identified based on the small gap between the ground truth and the predicted gaze trajectory.
Figure 7(b) highlights predictions related to Incorrect Execution. Here, our approach’s gaze predictions diverge from the ground truth, effectively flagging mistakes in action execution.